5. The sum of the page numbers on facing pages of a book is 157; what are the page numbers?
Answer by stanbon(75887)

The sum of the page numbers on facing pages of a book is 157; what are the page numbers?
-----------------
Those would be consecutive page numbers.
------------
Equation:
x + (x+1) = 157
2x = 156
x = 78 (1st page)
x+1 = 79 (next page)
==========================
Cheers,
Stan H.